14) The operations manager of a company that manufactures tiles wants to determine whether the quality of work is independent of the daily shifts. He randomly selects 384 tiles and carefully inspects them. Each tile is either classified as perfect, satisfactory, or defective, and the shift that produced it is also recorded. The data are summarised in the two-way table below and a ? ଶ test of independence is performed. Perfect Satisfactory Sempurna) Memuaskan 106 124 Defective (Rosak/ Total Jumlah/ 231 67 85 153 Shift 1 (Syif 11 Shift 2 (Syif

A) Find the critical value for the test at ? = 0.01.

A. 6.635 C. 10.597 B. 4.605 D. 9.210

Calculate the test statistics, ? ^2 to determine whether the level of quality is independent of the shifts.

A. 0.2354 C. 1.6452 B. 2.9513 D. 0.8662
Answer #1

Solution:

Column 1 Column 2 Column 3 Total Expected Values Row 1 173x231 384 = 104.07 209x231 384 = 125.727 2.231 = 1.203 231 384 Row 2

Squared Distances Column 1 Column 2 Column 3 Row 1 (106-104.07 104.07 = 0.036 (124-125.727) 125.727 0.024 (1-1.203)2 1.203 =